OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x + 7y = -21. x intercept = (3, 0); y intercept = (0, 21) x intercept = (-3, 0); y intercept = (0, -21) x intercept = (21, 0); y intercept = (0, 3) x intercept = (-21, 0); y intercept = (0, -3)

OpenStudy (anonymous):

For y intercept make x =0 0 + 7y = -21 solve. ( 0, ? ) for x intercept make y =0. x + 7(0)= -21 solve.( ? , 0 )

OpenStudy (anonymous):

-21/

OpenStudy (anonymous):

what is -21 ?

OpenStudy (anonymous):

give me the co-ordinates. (x , y)

OpenStudy (anonymous):

is x and y is 7?

OpenStudy (anonymous):

No solve y = -21/7.... what is y? keep working @primeralph will help you right its bed time.

OpenStudy (anonymous):

-3?

OpenStudy (primeralph):

@mebs You taunting me?

OpenStudy (anonymous):

i think the right answer is b

OpenStudy (primeralph):

At x = 0, y = -3 , (0,-3) at y = 0, x = -21, (-21,0)

OpenStudy (primeralph):

D.

OpenStudy (anonymous):

ugh. thxs!

OpenStudy (primeralph):

You're welcome.